NEAR-INFRARED DIODE LASER SPECTROSCOPY OF THE A–X SYSTEM OF N_{2}^{+} IN THE 1.2 - 1.3 μm REGION

\underline{C. TARSITANO} and T. OKA, Department of Chemistry, Department of Astronomy and Astrophysics, and the Enrico Fermi Institute, University of Chicago, Chicago, IL 60637.

Since Meinel’s identification during the auroral storm in 1950^{a}, the A^{2}Π_{u} ← X^{2}Σ_{g}^{+} Meinel system has been studied by a great number of spectroscopists in the wavelength region from 541 nm to 4.6 μm providing information from 27 bands. Two new bands, the (3,3) and (4,4) bands, of the Meinel system of N_{2}^{+} were observed using a near-infrared diode laser and the velocity modulation technique. A comprehensive analysis of the newly observed bands will be presented.

aA. Meinel, Astrophys. J. 112, 562 (1950) and 114, 431 (1951).